# Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

### v6

@@ 5.2. Finance
Applications
GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Cao, 2022). Key applications include customer-facing bots that provide interactive banking and wealth-management support, modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Cao, 2022), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Cao, 2022). Empirical deployments highlight these advances. For example, PayPal’s fraud detection reduced false positives by 50% in 2023 while processing 10 million transactions daily (Ngai et al., 2011). Charles Schwab’s assistant resolved 62% of client queries and improved satisfaction by 28% in 2023 (Wang et al., 2023b). JPMorgan’s simulator improved trading returns by 12% and reduced value-at-risk by 13% in 2023 (Zhang et al., 2023). HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025). Silva et al. also reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
ResGenAI challenges
@@
GenAI in finance raises a number of domain-specific risks that complicate its responsible adoption. These include discriminatory lending that replicates historical bias in credit and valuation data , security vulnerabilities where adversarial manipulation targets fraud detection or trading systems, consumer data privacy concerns arising from reliance on sensitive financial records, opacity of GenAI that limits auditability and accountability, hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice, and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (Hendry et al., 2022). Empirical use cases underscore these risks. For example, Zillow’s valuation model undervalued homes in minority neighborhoods and reduced appraised equity for underwriting (Adams et al., 2022). Citibank was also fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
Evaluation context
The primary finance risks map to C1 (lending discrimination), C3 (adversarial manipulation of trading/fraud systems), C6 (consumer data privacy), and C4 (hallucinated financial advice). No benchmark in Table 4 specifically targets lending discrimination or financial adversarial attacks. The most relevant KPIs are SPD and DI (KPIs 3–4) applied to loan approval rates across demographic groups, robustness $R$ (KPI 6) under financial adversarial scenarios, and audit frequency $F_{a}$ (KPI 10) aligned with SEC and OCC reporting cadences.
ResGenAI strategies
Multiple strategies are being prioritized to address these financial risks. Structured fairness audits (C1, KPIs 3–4) and fairness-aware data augmentation are increasingly required in lending and credit scoring. Post-hoc explainability (C9, KPIs 7–8) methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models. Privacy-preserving techniques (C6, KPI 2), including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data. Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets. Alignment with regulatory compliance (C10) frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022). Finally, embedding HITL oversight (C7, KPI 10) across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions.
5.3. Education
Applications
GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Holmes et al., 2022). Key applications include automated grading that accelerates evaluation and reduces instructor workload, personalized tutoring systems that adapt to student knowledge gaps and learning styles, content generation that produces adaptive curricula and study materials, accessibility tools that deliver multimodal and multilingual support for inclusive education, and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024). Recent deployments provide evidence of these advances. Carnegie Learning’s MATHia platform improved algebra test scores by 22% in U.S. high schools in 2023 (Carnegie Learning, 2023). A 2024 Arizona State University pilot found that students using an LLM-powered Socrates chatbot showed 42% higher engagement on surveys (Arizona State University, 2024). A Duolingo study in 2024 reported a 12% increase in learner retention and practice frequency through generative conversational exercises (Duolingo Research, 2024). A U.K. school district pilot in 2024 showed that 87% of students felt classroom chat assistants clarified concepts and 92% of teachers reported time saved (UK Department for Education, 2024).
ResGenAI challenges
Responsible use of GenAI in education faces several pressing concerns. These include cultural insensitivity where models fail to capture local norms and values, student data privacy and particularly minors’ data privacy given the sensitivity of educational records , skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving, and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Zawacki-Richter et al., 2019). Empirical use cases underscore these risks. For example, India’s government-run “Newton” homework bot trained only on U.S. and U.K. corpora recommended “pork stew” to rural students in Rajasthan in 2021, violating cultural and nutritional guidelines and prompting suspension of the tool for eleven weeks (UNESCO GEM Report, 2022). A 2023 randomized controlled trial at the University of Cyprus found that physics students allowed to use GPT-4 for weekly assignments scored twenty-three percent lower on conceptual-transfer exam questions despite higher weekly grades, demonstrating over-reliance effects (Pavlou et al., 2023). In 2022, the Utah State Board of Education cancelled a contract for an AI essay-scorer when auditors reported that the vendor could not document the feature vectors or training data, raising concerns about pedagogical opacity and compliance with state education law (Utah State Board of Education, 2022).
Evaluation context
Education risks concentrate on C6 (minors’ data privacy), C4 (hallucinated pedagogical content), C1 (cultural bias), and C7 (system-level failure in tutoring pipelines). No benchmark addresses C6 in the context of minors’ data, and the user inclusivity index $U$ (KPI 11) critical for accessibility in educational settings, lacks any standardized evaluation protocol. Skills-evasion effects (a form of C7 system-level failure) are entirely absent from current safety suites. Priority KPIs include $P$ (KPI 2) under COPPA/FERPA, $E_{h}$ (KPI 9) for pedagogically harmful outputs, and $U$ (KPI 11) for WCAG conformance.
ResGenAI strategies
Multiple strategies are being explored to ensure responsible generative AI in education. Fairness audits (C1, KPIs 3–4) and external ethics audits verify alignment with frameworks such as the UNESCO(UNESCO, 2021a) AI-ethics rubric and can keep bias metrics below accepted thresholds. Privacy-preserving techniques (C6, KPI 2) including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021). Clear regulation and policy guidelines (C10) developed by international and national bodies provide governance and accountability mechanisms. Finally, academic integrity tools (C5, C9) such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Zhang and Zhao, 2024).
5.4. Arts and Entertainment
Applications
GenAI is reshaping creative production and cultural participation in the arts and entertainment sector. Prominent applications include creative writing tools that assist in script, story, and lyric development, generative art and design platforms that expand aesthetic experimentation, music composition and production systems that automate melody, arrangement, and mastering, film production support for visual effects, dubbing, and storyboarding, personalized recommendation engines that curate media consumption at scale and virtual influencers and avatars that shape marketing and fan engagement (Crawford, 2021). Recent deployments provide evidence of these advances. For instance, Spotify’s Discover Weekly playlists improved user engagement by 40% in 2023, with subsequent adjustments in 2024 reducing genre bias (Anderson and Lee, 2023). A 2024 marketing campaign featuring AI-driven virtual influencers reached more than ten million users (Crawford, 2021; Traitter and Klein, 2022). Davis et al. released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
ResGenAI challenges
Responsible use of GenAI in arts and entertainment faces a series of ethical, legal, and cultural concerns. These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization, job displacement that threatens creative labor markets, cultural insensitivity when generative tools fail to respect community values or norms deepfakes and misinformation that erode authenticity and trust, authorship and agency disputes where creative ownership is contested, insufficient consent in data usage for training models, and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019). Empirical use cases underscore these risks. For example, the U.S. Copyright Office’s 2023 ruling in Thaler v. Perlmutter rejected copyright protection for AI-generated works, reinforcing the requirement of human authorship (U.S. Copyright Office, 2023). The public controversy around the 2023 song “Heart on My Sleeve,” which mimicked the voices of Drake and The Weeknd, raised critical questions about derivative works and copyright boundaries. In 2025, WeTransfer updated its terms of service to allow user content for AI training, sparking backlash over transparency and consent (WeTransfer, 2025).
Evaluation context
The dominant risks map to C5 (deepfakes and media integrity), C1 (representational bias in generated content), and C9 (attribution transparency). Only SHIELD and MM-SafetyBench provides explicit C5 evaluation in Table 4, and its scope does not extend to voice cloning or music mimicry, the vectors most salient for this sector. Copyright memorization, a C6-adjacent concern, is not addressed by any privacy benchmark. Priority KPIs are currently underspecified for creative domains; provenance-verification rates and consent-compliance metrics (analogous to KPI 2) are needed but lack standardized protocols.
ResGenAI strategies
Multiple strategies are being adopted to promote responsible use of generative AI in arts and entertainment. Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value. Structured bias audits (C1, KPIs 3–4) and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping . Deepfake detection tools (C5) are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019). Collective bargaining and union agreements such as those negotiated by SAG-AFTRA establish protections for creative workers (SAG-AFTRA, 2023). International ethical guidelines (C10) and transparency labeling requirements provide standards for disclosure and attribution. Finally, intellectual property management systems (C6, C9) are being advanced to secure consent and compliance. For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these systems require explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.
5.5. Agriculture
Applications
GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation. Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage, crop yield prediction models that integrate weather, soil, and satellite data, pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy, virtual agronomy advisors that deliver context-sensitive recommendations to farmers, agricultural robotics for automated harvesting and monitoring, crop breeding simulations that accelerate trait discovery, and weather forecasting for climate adaptation (Costa et al., 2023). While platforms like Microsoft’s FarmBeats (Microsoft Research, 2023) initially achieved significant yield and water-saving gains through IoT and predictive ML , GenAI is now being integrated to augment these systems through generative weather simulation for risk modeling and as LLM-based natural language interfaces . These tools leverage GenAI to translate complex sensor data into actionable, localized advice.
ResGenAI challenges
Responsible adoption of GenAI in agriculture faces several interrelated challenges. These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions, farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms, bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers, overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom, IoT security vulnerabilities that expose connected farm systems to cyberattacks, and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; Sapkota et al., 2024). Empirical use cases underscore these risks. For example, farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge, illustrating both hallucination and cultural misalignment. Cases of contested ownership emerged when farm sensor data collected by multinational seed companies were commercialized without equitable farmer compensation (Bronson, 2019). IoT breaches in 2023 European trials highlighted vulnerabilities in sensor networks that exposed proprietary farm-level data.
Evaluation context
Agriculture risks primarily concern C4 (hallucinated agronomic advice), C6 (farm data ownership and privacy), and C7 (system-level failure in IoT-integrated advisory pipelines). No benchmark in our assessment provides geo-aware hallucination testing, and C7 remains untested for sensor-to-advice chains. Priority KPIs are $E_{h}$ (KPI 9) operationalized as harmful agronomic recommendations per $N$ advisory sessions, $P$ (KPI 2) for data governance compliance, and $E_{\text{kWh}}$ (KPI 12) for sustainability reporting—particularly relevant given the energy constraints of low-resource deployment contexts.
RAI strategies
Multiple strategies are being pursued to enable responsible deployment of generative AI in agriculture. Geo-aware retrieval and annotation (C4, KPI 9) techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks. Data governance contracts (C6, KPI 2) are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets. Embedding human-in-the-loop oversight (C7, KPI 10) into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation. Privacy-preserving IoT integration (C6, KPI 2), including differential privacy and federated learning, protects sensitive farm data across institutions. Finally, alignment with sustainability regulations (C10, KPI 12) such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards. Together, these approaches enhance both the equity and safety of generative AI in global agriculture.
@@ 5.6. Defense
Applications
GenAI is increasingly influencing defense through simulation, sensing, and decision support. Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development, threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness, autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments, and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (News, 2024). Recent deployments provide evidence of these advances. For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018). The U.S. Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (News, 2024). NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis.
RAI challenges
Responsible adoption of GenAI in defense introduces ethical operational and legal concerns. These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved, opacity of decision making that limits auditability and traceability in high-stakes contexts, adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning, escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk, ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement, research security concerns over dual-use model capabilities and insider threat, protection of critical infrastructure as AI is integrated into command networks and sensors, and facial recognition deployment where demographic performance differences raise risks in security operations. Empirical use cases underscore these risks. For example, national and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints that required added human oversight (NATO Allied Command Transformation, 2024). Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defences if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025). Facial recognition evaluations documented demographic differentials that necessitate strict control and testing before operational use (Grother et al., 2019, 2022).
Evaluation context
Defense risks concentrate on C3 (adversarial robustness), C7 (system-level failure in multi-agent autonomy), C8 (red-team realism), and C10 (governance alignment with IHL and NATO principles). C8 is the most critical gap: only 2 of 21 benchmarks (Rainbow Teaming and ALERT) score above zero for adaptive adversary simulation, yet defense applications face the most sophisticated threat actors. C7 is entirely absent for multi-agent autonomous systems. Priority KPIs are robustness $R$ (KPI 6) under military-grade adversarial scenarios, $E_{h}$ (KPI 9) for consequential targeting or escalation errors, and audit frequency $F_{a}$ (KPI 10) aligned with NATO after-action review cadences.
RAI strategies
Multiple strategies are being advanced to ensure responsible deployment. Human in the loop oversight (C7, C8, KPI 10) is prioritized for mission-critical tasks so that trained operators validate model outputs before action. Meaningful human control (C10) is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law. Explainable AI for defense transparency (C9, KPIs 7–8) adopts traceability and interpretability requirements to support audit and after-action review. Red teaming (C3, C8, KPI 6) and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding. International cooperation (C10) aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024). Ethical review boards and program-level governance embed policy risk assessment data governance and model documentation across the AI life cycle (U.S. Department of Defense, 2024; Defense Innovation Unit, 2020).
5.7. Cross-Applications Synthesis of GenAI Maturity
@@
To provide a critical comparison of how responsible practices are being operationalized across domains, and to connect sectoral risks to the evaluation infrastructure in Sections 4.24.4, we present the extended Synthesis Matrix in Table 9. Beyond risk mechanisms and regulatory drivers, the table now maps each sector to its primary rubric-criteria gaps and priority lifecycle KPIs, making explicit which evaluation evidence is available and where portable audit artifacts are missing.
5.7.1. Stakeholder-Metric Alignment: Operationalizing the Framework
@@
To ensure the sectoral insights in Section 5 translate into actionable governance, we map the relationship between stakeholder roles and our evaluation infrastructure. Based on our synthesized findings, technical frameworks often fail because they do not specify which evidence is required by which actor. In our model, Model Developers utilize the robustness (C3) and toxicity (C2) scores from Table 4 to guide iterative alignment during the fine tuning phase. Conversely, Compliance Auditors and Regulators prioritize the statistical parity (SPD/DI) KPIs (Section 4.2.4) and governance-alignment criteria (C10) to verify adherence to mandates like the EU AI Act (Outeda, 2024) or NIST AI RMF (of Standards and, NIST). Finally, Domain Practitioners in high-stakes fields like Healthcare and Finance rely on the High-Stakes Error Rate ($E_{h}$) and local faithfulness gaps (LFG) to determine the necessary level of human-in-the-loop oversight. By bridging domain-specific mechanisms with normalized technical metrics, our framework provides the ”chain of trust” required to move ResGenAI from theoretical safety to verifiable institutional accountability.
Table 9. Synthesis Matrix: Comparative Analysis of RAI Maturity, Regulatory Pressures, Rubric Criteria Gaps, and Priority KPIs Across Domains.
Sector Dominant Risk Mechanism Technical Control Maturity Primary Regulatory Driver Primary C1 to C10 Gaps Priority KPIs
Healthcare Medical hallucination (Lyu et al., 2025); Diagnostic bias (Gonzalez et al., 2022) Medium: RAG (Wu et al., 2024), HITL, Diff. Privacy (Kaissis et al., 2021) FDA Guidance (U.S. Food and Drug Administration, 2024); HIPAA (Shiraz and Lee, 2019) C4: no clinical hallucination benchmark; C6: no HIPAA specific privacy suite; C7: 4/21 cover system failure $E_{h}$ (9), $P$ (2), SPD/DI (3 to 4), $\Delta_{\text{acc}}$ (5)
Finance Lending discrimination (Mehrabi et al., 2021b); Market manipulation (Hendry et al., 2022) High: Fairness audits (Pfohl et al., 2022); Multi party comp. (Khalid et al., 2023) SEC (U.S. Securities and Exchange Commission, 2021); OCC (Office of the Comptroller of the Currency, 2020) C1: no lending discrimination benchmark; C3: no financial adversarial attack suite SPD/DI (3 to 4), $R$ (6), $F_{a}$ (10)
Education Student privacy (Livingstone et al., 2021); Skills evasion (Pavlou et al., 2023) Low: Watermarking (Jiang et al., 2023); Fed. learning (Yang, 2021) UNESCO AI Rubric (UNESCO, 2021b); State Edu. Laws (Utah State Board of Education, 2022) C6: no minors data protocol; C7: skills evasion untested; KPI 11 lacks standard $P$ (2), $E_{h}$ (9), $U$ (11)
[…]
Agriculture Ownership conflicts (Bronson, 2019); Global South exclusion (Tzachor et al., 2022) Low: Geo aware RAG (Ma et al., 2023); Data contracts (Bronson, 2019) FAO (Costa et al., 2023); EU CAP (Costa et al., 2023) C4: no geo aware hallucination test; C7: IoT advisory pipelines untested $E_{h}$ (9), $P$ (2), $E_{\text{kwh}}$ (12)
Defense Autonomous lethality (International Committee of the Red Cross, 2021); Adversarial spoofing (National Institute of Standards and Technology, 2025a) High: Red teaming (Corporation, 2024); Meaningful HITL (International Committee of the Red Cross, 2021) NATO Principles (NATO, 2024); IHL (United Nations Institute for Disarmament Research, 2025) C8: only 2/21 benchmarks score >>0; C7: absent for multi agent autonomy $R$ (6), $E_{h}$ (9), $F_{a}$ (10)
Legal & Justice Recidivism prediction bias (Skeem and Lowenkamp, 2016); Automated decision opacity (Chesterman, 2021) Low to Med.: Explainability (LIME/SHAP) (Microsoft Research, [n. d.]); Bias audits (Center for Data Science and Public Policy, [n. d.]) EU AI Act Art. 6 (Edwards, 2021); U.S. DOJ Guidelines (Podgor, 2003) C1: no recidivism specific fairness benchmark; C2: legal advice toxicity untested; C10: judicial HITL criteria absent SPD/DI (3 to 4), $E_{h}$ (9), $R$ (6), $F_{a}$ (10)
Manufacturing Predictive maintenance failures (Hashemian, 2010); Adversarial sensor spoofing (Cao et al., 2019) Medium: Digital twins (Mihai et al., 2022); Anomaly detection (Chandola et al., 2009) IEC 62443 (ISO/IEC, 2021b); EU Machinery Reg. (European Union, 2024) C3: no industrial adversarial robustness suite; C4: no sensor drift benchmark; C7: multi robot coordination untested $E_{h}$ (9), $R$ (6), $E_{\text{kWh}}$ (12), $F_{a}$ (10)

### v7

@@ 5.2. Finance
Applications and challenges. GenAI is transforming fraud detection, credit scoring, risk forecasting, and customer-facing advisory. Domain-specific risks include discriminatory lending that replicates historical bias in credit and valuation data, adversarial manipulation of trading and fraud systems, consumer data privacy across jurisdictions, and hallucinated financial advice in customer-facing systems. Zillow’s valuation model undervalued homes in minority neighborhoods, reducing appraised equity for underwriting (Adams et al., 2022). Citibank was fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
Domain-specific strategies. Structured fairness audits in lending must go beyond standard SPD/DI to address intersectional and proxy-variable discrimination embedded in credit histories. Stronger data provenance frameworks are needed to track the lineage and validity of synthetic financial datasets. Robustness testing (C3, KPI 6) against financial adversarial scenarios (e.g., adversarial inputs to trading models) is sector-critical but absent from current benchmarks in Table 3. Audit frequency $F_{a}$ (KPI 10) should align with SEC and OCC reporting cadences, and secure multiparty computation enables cross-institutional risk modelling without exposing proprietary data.
@@ 5.3. Defense
Applications and challenges. GenAI is increasingly used for simulation and war-gaming, multi-INT threat detection, autonomous reconnaissance, and logistics optimization. Domain-specific risks include autonomous weapons and responsibility gaps, opacity of decision-making in high-stakes targeting, adversarial vulnerabilities (spoofing, data poisoning), escalation of conflicts through compressed decision windows, and dual-use research security. National and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints requiring additional human oversight (NATO Allied Command Transformation, 2024).
Domain-specific strategies. Defense is the domain where red teaming (C3, C8, KPI 6) is most critical, yet only Rainbow Teaming and ALERT score above zero for adaptive adversary simulation in our assessment. Meaningful human control (C10) must keep decisions on the use of force within accountable human authority, consistent with international humanitarian law and NATO’s Principles of Responsible Use (NATO, 2024). Explainability requirements (C9) in defense emphasize traceability for after-action review rather than end-user understanding. System-level failure testing (C7) for multi-agent autonomous systems is entirely absent from current safety suites, representing the single largest benchmark gap for this sector. Ethical review boards and program-level governance embed policy, risk assessment, data governance, and model documentation across the AI lifecycle (U.S. Department of Defense, 2024).
5.4. Other Domains: Education, Arts, and Agriculture
Education. GenAI supports personalized tutoring, automated grading, and adaptive content generation, but responsible use is complicated by minors’ data privacy (C6), cultural insensitivity (C1), skills evasion where over-reliance on AI weakens critical thinking (C7), and pedagogical opacity. A 2023 randomized controlled trial at the University of Cyprus found that physics students using GPT-4 for assignments scored 23% lower on conceptual-transfer exam questions despite higher weekly grades (Pavlou and others, 2023). No current benchmark addresses C6 for minors’ data, and skills-evasion effects (a form of system-level failure) are entirely absent from safety suites. Domain-specific needs include compliance with COPPA/FERPA (KPI 2), academic integrity tools (C5, C9) such as watermarking and output tracking, and accessibility metrics aligned with WCAG (user inclusivity index $U$, KPI 11).
Arts and entertainment. GenAI is reshaping creative production but raises concerns about copyright infringement (C6), deepfakes (C5), cultural insensitivity (C1), and contested authorship and attribution (C9). The U.S. Copyright Office’s 2023 ruling in Thaler v. Perlmutter rejected copyright protection for AI-generated works (U.S. Copyright Office, 2023). Only SHIELD provides explicit deepfake evaluation in Table 3, and its scope excludes voice cloning and music mimicry. Domain-specific strategies include artist collaboration platforms and alternative compensation models, intellectual property management systems requiring explicit consent for digital replicas, and collective bargaining agreements such as SAG-AFTRA protections.
Agriculture. GenAI supports precision farming, pest detection, and advisory chatbots, but responsible adoption faces hallucinated agronomic advice (C4), farm data ownership conflicts (C6), bias toward major corporations and regions (C1), and IoT security vulnerabilities (C3). Farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge. No benchmark provides geo-aware hallucination testing or sensor-to-advice chain validation (C7). Domain-specific needs include geo-aware retrieval and annotation to contextualize outputs with local conditions, data governance contracts clarifying ownership and revenue-sharing for farm-generated datasets, and sustainability reporting (KPI 12) aligned with FAO and EU Common Agricultural Policy frameworks.
5.5. Cross-Domain Synthesis of GenAI Maturity
@@
To provide a critical comparison of how responsible practices are being operationalized across domains, and to connect sectoral risks to the evaluation infrastructure in Sections 4.24.4, we present the extended Synthesis Matrix in Appendix Table 6 that maps each sector to its primary rubric-criteria gaps and priority lifecycle KPIs.
@@
Based on our synthesized findings, technical frameworks often fail because they do not specify which evidence is required by which actor. In our model, Model Developers utilize the robustness (C3) and toxicity (C2) scores from Table 3 to guide iterative alignment during fine-tuning. Compliance Auditors and Regulators prioritize the SPD/DI KPIs (Section 4.2.4) and governance-alignment criteria (C10) to verify adherence to mandates like the EU AI Act (Outeda, 2024) or NIST AI RMF (of Standards and (NIST), 2024). Domain Practitioners in high-stakes fields like Healthcare and Finance rely on the high-stakes error rate ($E_{h}$) and local faithfulness gaps (LFG) to determine the necessary level of HITL oversight.